# UniFashion: A Unified Vision-Language Model for Multimodal Fashion Retrieval and Generation

### v1

@@ Abstract
The fashion domain encompasses a variety of real-world multimodal tasks, including multimodal retrieval and multimodal generation. The rapid advancements in artificial intelligence generated content, particularly in technologies like large language models for text generation and diffusion models for visual generation, have sparked widespread research interest in applying these multimodal models in the fashion domain. However, tasks involving embeddings, such as image-to-text or text-to-image retrieval, have been largely overlooked from this perspective due to the diverse nature of the multimodal fashion domain. And current research on multi-task single models lack focus on image generation. In this work, we present UniFashion, a unified framework that simultaneously tackles the challenges of multimodal generation and retrieval tasks within the fashion domain, integrating image generation with retrieval tasks and text generation tasks. UniFashion unifies embedding and generative tasks by integrating a diffusion model and LLM, enabling controllable and high-fidelity generation. Our model significantly outperforms previous single-task state-of-the-art models across diverse fashion tasks, and can be readily adapted to manage complex vision-language tasks. This work demonstrates the potential learning synergy between multimodal generation and retrieval, offering a promising direction for future research in the fashion domain. The source code is available at https://github.com/xiangyu-mm/UniFashion.
@@
Xiangyu Zhao, Yuehan Zhang, Wenlong Zhang, Xiao-Ming Wu Department of Computing, The Hong Kong Polytechnic University {xiang-yu.zhao, wenlong.zhang}@connect.polyu.hk, john_love@whu.edu.cn, xiao-ming.wu@polyu.edu.hk
@@ 1 Introduction
The fashion domain presents a range of real-world multimodal tasks, encompassing multimodal retrieval (Gao et al., 2020; Wu et al., 2021; Bai et al., 2023) and multimodal generation (Yang et al., 2020) tasks. Such tasks have been utilized in diverse e-commerce scenarios to enhance product discoverability, seller-buyer interaction, and customer conversion rates after catalog browsing (Han et al., 2023; Zhuge et al., 2021). The remarkable progress in the field of artificial intelligence generated content (AIGC), particularly in technologies like large language models (LLMs) (Chiang et al., 2023; Touvron et al., 2023; Brown et al., 2020) for text generation and diffusion models Rombach et al. (2022); Nichol et al. (2022); Saharia et al. (2022) for visual generation, has sparked widespread research interest in applying these multimodal models in the fashion domain.
Multimodal large language models (Liu et al., 2023a; Dai et al., 2023; Dong et al., 2023) (MLLMs) seem to emerge as a promising direction for a single multi-task model. However, due to the heterogeneous nature of the multimodal fashion tasks (Han et al., 2023), existing MLLMs lack the capability to be directly applied to the fashion domain, such as embedding ability. For example, in the fashion domain, retrieval tasks that rely on embedding ability, like image-to-text or text-to-image retrieval, have been largely neglected from this aspect. Furthermore, MLLMs lack the ability to solve composed image retrieval (CIR) (Liu et al., 2021; Baldrati et al., 2022) task, which composes the reference image and related caption into a joint embedding to calculate similarities with the candidate images and is particularly relevant in fashion recommendation systems (Han et al., 2017).
Drawing inspiration from GRIT (Muennighoff et al., 2024), which successfully combined embedding and generative tasks into a unified model for text-centric applications and showed improved embedding performance through the addition of a generative objective, it becomes clear that investigating task correlations and integrating embedding with generative models in the fashion domain is is both necessary and promising.
@@
While previous works (Han et al., 2023; Zhuge et al., 2021) in the fashion domain have also proposed using a single model for solving multiple tasks, they ignore the image generation tasks. Besides, for fashion tasks such as try-on (Choi et al., 2021) and fashion design (Baldrati et al., 2023b), it is generally required to generate target images based on multimodal input. However, previous works (Baldrati et al., 2023b) in fashion image generation typically adopt the CLIP text encoder to encode text information, which may not be capable of effectively understanding the textual context due to their weaker text encoder, as noted in Saharia et al. (2022). Hence, we posit that current studies have not fully exploited the potential in learning synergy between generation and retrieval.
@@
Secondly, our model enhances performance via mutual task reinforcement. Specifically, the caption generation module aids the CIR task, while jointly training the generation and retrieval tasks improves the multimodal encoder for the diffusion module.
Third, extensive experiments on diverse fashion tasks—including cross-modal retrieval, composed image retrieval, and multimodal generation—demonstrate that our unified model significantly surpasses previous state-of-the-art methods.
@@ 2.1 Fashion Tasks
Fashion tasks encompass a range of image and language manipulations, including cross-modal retrieval, composed image retrieval, fashion image captioning and generation, etc. The representative tasks can be briefly divided into the following two groups:
Fashion Retrieval generally consists of Cross-Modal Retrieval (CMR) (Ma et al., 2022; Rostamzadeh et al., 2018) and composed image retrieval (CIR) tasks (Baldrati et al., 2023a; Bai et al., 2023). CMR requests to efficiently retrieve the most matched image/sentence from a large candidate pool $\mathcal{D}$ given a text/image query. CIR is a special type of image retrieval with a multimodal query (a combination of a reference image and a modifying text) matched against a set of images. It retrieves a target image from a vast image database based on a reference image and a text description detailing changes to be applied to the reference image. In this scenario, a query pair $p=\{I_{R},t\}$ is provided, where $I_{R}$ is the reference image and $t$ is the text describing the desired modifications. The challenge for this task is to accurately identify the target image $I_{T}$ that best matches the query among all potential candidates in the image corpus $\mathcal{D}$.
Fashion Generation consists of Fashion Image Captioning (FIC) and Fashion Image Generation (FIG). FIC (Yang et al., 2020) aims to generate a descriptive caption for a product based on the visual and/or textual information provided in the input. FIG aims to generate images based on the multimodal input, such as try-on (Choi et al., 2021; Gou et al., 2023) and fashion design (Baldrati et al., 2023b).
@@ 2.2 Multimodal Language Models
Recent research has witnessed a surge of interest in multimodal LLMs, including collaborative models (Wu et al., 2023; Yang et al., 2023b; Shen et al., 2023) and end-to-end methods (Alayrac et al., 2022; Zhao et al., 2023; Li et al., 2022; Bao et al., 2021; Wang et al., 2022b, a, a). More recently, some works also explore training LLMs with parameter-efficient tuning (Li et al., 2023b; Zhang et al., 2023b) and instruction tuning (Dai et al., 2023; Liu et al., 2023a; Ye et al., 2023; Zhu et al., 2023a; Li et al., 2023a). They only focus on generation tasks, while UniFashion is built upon a unified framework that enables both retrieval and generation tasks.
@@ 3.1 Phase 1: Cross-modal Pre-training
In the first stage, we conduct pre-training on the retrieval and generation modules to equip the Large Language Model (LLM) and diffusion model with strong cross-modal fashion representation capabilities for the next phase.
@@ 3.2.2 Composed Multimodal Generation
Table 2: Image captioning task performance on the FashionGen dataset.
@@
Table 3: Performance analysis of unpaired settings on VITON-HD and MGD datasets across different input modalities.
@@ 4.1 Experimental Setup
We initialize the multimodal encoder from BLIP2’s Q-Former and MLLM from LLaVA-1.5. As for the diffusion module, following, StableVITON, we inherit the autoencoder and the denoising U-Net of the Stable Diffusion v1.4. We initialize the weights of the U-Net from the Paint-by-Example and for more refined person texture, we utilized a VAE fine-tuned on the VITONHD dataset from StableVITON. The statistics of the two-stage datasets can be found in Table 6. For cross-modal retrieval, we evaluated UniFashion on FashionGen validation set. For the image captioning task, UniFashion is evaluated in the FashionGen dataset. For the composed image retrieval task, we evaluated the Fashion-IQ validation set. To maintain consistency with previous work, for the composed image generation task, we fine-tuned UniFashion and evaluated it on the VITON-HD and MGD datasets. More details can be found in Appendix C.
Phase 1: For multimodal representation learning, we follow BLIP2 and pretrain the Q-Former on fashion image-text pairs. To adapt the model for multimodal generation, we freeze the parameters of Q-Former and fine-tune the MLLM and diffusion model with their task specific adapters separately. Due to the different styles of captions in different fashion datasets, we adopt the approach of instruction tuning to train the LLM so that it can generate captions of different styles. More details can be found in Appendix D.
Phase 2: In order to make UniFashion have the composed retrieval and generation abilities, we freeze the parameters of LLM and diffusion model, only fine-tune the multimodal encoder.
@@ 4.2 Evaluation Methods
Cross-modal retrieval evaluation: We consider both image-to-text retrieval and text-to-image retrieval with random 100 protocols used by previous methods. 100 candidates are randomly sampled from the same category to construct a retrieval database. The goal is to locate the positive match depicting the same garment instance from these 100 same-category negative matches. We utilize Recall@K as the evaluation metric, which reflects the percentage of queries whose true target ranked within the top K candidates.
Fashion image captioning evaluation: For evaluating the performance of caption generation, we utilize BLEU-4, METEOR, ROUGE-L, and CIDEr as metrics.
Composed fashion image retrieval evaluation: We compare our UniFashion with CIR methods and the FAME-ViL model of V + L that is oriented towards fashion in the original protocol used by Fashion-IQ. For this task, we also utilize Recall@K as the evaluation metric.
Composed fashion image generation evaluation: We compare our UniFashion with try-on methods on VITON-HD dataset and fashion design works on MGD dataset. To evaluate the quality of image generation, we use the Frechet Inception Distance (FID) score to measure the divergence between two multivariate normal distributions and employ the CLIP Score (CLIP-S) provided in the TorchMetrics library to assess the adherence of the image to the textual conditioning input (for fashion design task).
@@ 4.3 Comparative Analysis of Baselines and Our Method
UniFashion performs better compared to baselines in all data sets. Tab. 1 presents the evaluation results for each baseline and our models in FashionGen data sets for cross-modal retrieval. UniFashion outperforms most of the baseline models on both the text-to-image and image-to-text tasks. Following FAME-ViL, we also adopt a more challenging and practical protocol that conducts retrieval on the entire product set, which is in line with actual product retrieval scenarios. In Tab. 2, we performed a comparison between our UniFashion and other baselines on the FashionGen dataset for the image captioning task. By integrating the powerful generative ability of the LLM, our model performed significantly better than the traditional multimodal models in this task. In Tab. 4, we conducted a comparison between our UniFashion and CIR-specialist methods. Our findings are in line with those of Tab. 1.
After fine-tuning UniFashion on different modal input composed image generation/editing tasks, it also demonstrates excellent performance. Tab. 3 evaluates the quality of the generated image of UniFashion in the VITON-HD unpaired setting. In order to verify that our model can achieve good results in a variety of modal inputs, we have conducted tests respectively on the traditional try-on task and the fashion design task proposed in MGD. For a fair evaluation with baselines, all the models are trained at a 512 × 384 resolution. To confirm the efficacy of our approach, we assess the realism using FID and KID score on all the tasks and using CLIP-S score for fashion design task. As can be seen, the proposed UniFashion model consistently outperforms competitors in terms of realism (i.e., FID and KID) and coherence with input modalities (i.e., CLIP-S), indicating that our method can better encode multimodal information. Meanwhile, although our model is slightly lower than StableVITON on the try-on task, this is because we froze the parameters of the diffusion model on the try-on task and only fine-tuned the Q-former part, but it can still achieve top2 results.
@@ 4.4 Ablation Study
Our model completes the multimodal composed tasks in more aspects. In Tab. 4, we also carry out ablation studies on different retrieval methods. Since UniFashion is capable of generating captions, for the CIR task, we initially utilize UniFashion to generate the captions of candidate images and then conduct the image retrieval task (denoted as UniFashion w/o cap) and the caption retrieval task (denoted as UniFashion w/o img). We find that our single-task variant has already achieved superior performance in the relevant field. Furthermore, due to the generative ability of our model, the pregenerated candidate library optimizes the model’s performance in this task. For specific implementation details, please refer to Appendix B.
We researched the impact of different modules in UniFashion on various fashion tasks. In Tab. 5, we perform an ablation study on the proposed model architecture, with a focus on LLM and diffusion models. For comparison on the cross-modal retrieval task (CMR), we design the base model as directly fine-tuning BLIP2 without any new modules. The results indicate that the base model performs relatively well on this task and that the introduction of other modules does not lead to significant improvements. However, in the CIR task, the introduction of LLM and diffusion models as supervision can lead to significant improvements, especially when utilizing the captions pregenerated by the UniFashion to assist in retrieval, resulting in greater benefits. At the same time, we note that, after introducing the diffusion model, it may have some negative impact on the model’s image captioning ability, possibly due to the inherent alignment differences between LLM and the diffusion model.

### v2

@@ Abstract
The fashion domain includes a range of real-world multimodal tasks, such as multimodal retrieval and generation. Recent advancements in AI-generated content, particularly large language models for text and diffusion models for visuals, have spurred significant research interest in applying these multimodal models to fashion. However, fashion models must also effectively handle embedding tasks, like image-to-text and text-to-image retrieval. Moreover, current unified fashion models often lack the capability for image generation. In this work, we present UniFashion, a unified framework that tackles the challenges of multimodal generation and retrieval tasks in the fashion domain, by integrating image and text generation with retrieval tasks. UniFashion unifies embedding and generative processes through the use of a diffusion model and LLM, enabling controllable and high-fidelity generation. Our model significantly outperforms previous state-of-the-art models focused on single tasks across various fashion-related challenges and can be easily adapted to manage complex vision-language tasks. This study highlights the synergistic potential between multimodal generation and retrieval, offering a promising avenue for future research in the fashion domain. The source code is available at https://github.com/xiangyu-mm/UniFashion.
@@
Xiangyu Zhao1, Yuehan Zhang2, Wenlong Zhang1,3, Xiao-Ming Wu1 1Department of Computing, The Hong Kong Polytechnic University 2Wuhan University, 3Shanghai AI Laboratory xiang-yu.zhao@connect.polyu.hk, xiao-ming.wu@polyu.edu.hk
@@ 1 Introduction
The fashion domain presents a range of real-world multimodal tasks, encompassing multimodal retrieval (Gao et al., 2020; Wu et al., 2021; Bai et al., 2023; Liu et al., 2024b) and multimodal generation (Yang et al., 2020) tasks. Such tasks have been utilized in diverse e-commerce scenarios to enhance product discoverability, seller-buyer interaction, and customer conversion rates after catalog browsing (Han et al., 2023; Zhuge et al., 2021). The remarkable progress in the field of artificial intelligence generated content (AIGC), particularly in technologies like large language models (LLMs) (Chiang et al., 2023; Touvron et al., 2023; Brown et al., 2020) for text generation and diffusion models Rombach et al. (2022); Nichol et al. (2022); Saharia et al. (2022) for visual generation, yielding significant advancements in numerous downstream tasks (Feng et al., 2023; Zhang et al., 2022) and sparking widespread research interest in applying these multimodal models to the fashion domain.
Instruction-tuned multimodal large language models (Liu et al., 2023a; Dai et al., 2023; Dong et al., 2023; Zhao et al., 2024) (MLLMs) have emerged as a promising direction for developing a single multi-task model (Shi et al., 2023). However, due to the heterogeneous nature of multimodal fashion tasks (Han et al., 2023), most existing MLLMs struggle to be directly applicable in the fashion domain. For example, in the fashion domain, retrieval tasks that rely on embedding ability, such as image-to-text or text-to-image retrieval, have largely been overlooked. Furthermore, existing MLLMs lack the ability to solve the composed image retrieval (CIR) (Liu et al., 2021; Baldrati et al., 2022) task, which composes the reference image and related caption in a joint embedding to calculate similarities with candidate images and is particularly relevant in recommender systems (Han et al., 2017; Liu et al., 2022, 2024a).
Drawing inspiration from GRIT (Muennighoff et al., 2024), which successfully combined generative and embedding tasks into a unified model for text-centric applications and enhanced embedding performance by incorporating a generative objective, it is evident that exploring task correlations and integrating embedding with generative models in the fashion domain is promising.
@@
While previous works (Han et al., 2023; Zhuge et al., 2021) in the fashion domain have also proposed using a single model for solving multiple tasks, they ignore image generation tasks. Besides, for fashion tasks such as try-on (Choi et al., 2021) and fashion design (Baldrati et al., 2023b), it is generally required to generate target images based on multimodal input. However, previous works (Baldrati et al., 2023b) in fashion image generation typically adopt the CLIP text encoder for encoding text information. This approach may not effectively capture the textual context due to the limitations of the text encoder, as noted by Saharia et al. (2022). Hence, we posit that current studies have yet to fully explore the potential synergy between generation and retrieval.
@@
  • Secondly, our model enhances performance via mutual task reinforcement. Specifically, the caption generative module aids the CIR task, while jointly training the generation and retrieval tasks improves the multimodal encoder for the diffusion module.
  • Thirdly, extensive experiments on diverse fashion tasks—including cross-modal retrieval, composed image retrieval, and multimodal generation—demonstrate that our unified model significantly surpasses previous state-of-the-art methods.
@@ 2.1 Fashion Tasks
Fashion tasks encompass a range of image and language manipulations, including cross-modal retrieval, composed image retrieval, fashion image captioning and generation, etc. The representative tasks can be briefly divided into the following two groups.
Fashion Retrieval.
It generally consists of Cross-Modal Retrieval (CMR) (Ma et al., 2022; Rostamzadeh et al., 2018) and composed image retrieval (CIR) tasks (Baldrati et al., 2023a; Bai et al., 2023). CMR requests to efficiently retrieve the most matched image/sentence from a large candidate pool $\mathcal{D}$ given a text/image query. CIR is a special type of image retrieval with a multimodal query (a combination of a reference image and a modifying text) matched against a set of images. It retrieves a target image from a vast image database based on a reference image and a text description detailing changes to be applied to the reference image. In this scenario, a query pair $p=\{I_{R},t\}$ is provided, where $I_{R}$ is the reference image and $t$ is the text describing the desired modifications. The challenge for this task is to accurately identify the target image $I_{T}$ that best matches the query among all potential candidates in the image corpus $\mathcal{D}$.
Fashion Generation.
It consists of Fashion Image Captioning (FIC) and Fashion Image Generation (FIG). FIC (Yang et al., 2020) aims to generate a descriptive caption for a product based on the visual and/or textual information provided in the input. FIG aims to generate images based on the multimodal input, such as try-on (Choi et al., 2021; Gou et al., 2023) and fashion design (Baldrati et al., 2023b).
@@ 2.2 Multimodal Language Models
Recent research has witnessed a surge of interest in multimodal LLMs, including collaborative models (Wu et al., 2023; Yang et al., 2023b; Shen et al., 2023) and end-to-end methods (Alayrac et al., 2022; Zhao et al., 2024; Li et al., 2022; Bao et al., 2021; Wang et al., 2022b, a, a). More recently, some works also explore training LLMs with parameter-efficient tuning (Li et al., 2023b; Zhang et al., 2023b) and instruction tuning (Dai et al., 2023; Liu et al., 2023a; Ye et al., 2023; Zhu et al., 2023a; Li et al., 2023a). They only focus on generation tasks, while our model UniFashion is designed as a unified framework that enables both retrieval and generation tasks.
@@ 3.1 Phase 1: Cross-modal Pre-training
In the first stage, we conduct pre-training on the retrieval and generative modules to equip the Large Language Model (LLM) and diffusion model with strong cross-modal fashion representation capabilities for the next phase.
@@ 3.2.2 Composed Multimodal Generation
3.3 Instruction-Tuning LLMs for Different Caption Style
USER: <Img><queries></Img> + Instruction. Assistant: <answer>.
@@ 3.3 Instruction-Tuning LLMs for Different Caption Style
Table 2: The Performance of UniFashion in the image captioning task on the FashionGen dataset.
@@
Table 3: Performance analysis of unpaired settings on the VITON-HD and MGD datasets across different input modalities.
@@ 4.1 Experimental Setup
We initialize the multimodal encoder using BLIP2’s Q-Former. Following the approach of LLaVA-1.5 (Liu et al., 2023a), we initialize the LLM from Vicuna-1.5 (Zheng et al., 2023). For the diffusion module, we adopt the autoencoder and denoising U-Net from Stable Diffusion v1.4, as utilized in StableVITON. The weights of the U-Net are initialized from Paint-by-Example. To achieve more refined person textures, we employ a VAE that has been fine-tuned on the VITONHD dataset, as done in StableVITON. The statistics of the two-stage datasets can be found in Table 6. For cross-modal retrieval, we evaluated UniFashion on FashionGen validation set. For the image captioning task, UniFashion is evaluated in the FashionGen dataset. For the composed image retrieval task, we evaluated the Fashion-IQ validation set. To maintain consistency with previous work, for the composed image generation task, we fine-tuned UniFashion and evaluated it on the VITON-HD and MGD datasets. More details can be found in Appendix B.
Phase 1:
For multimodal representation learning, we follow BLIP2 and pretrain the Q-Former on fashion image-text pairs. To adapt the model for multimodal generation, we freeze the parameters of Q-Former and fine-tune the MLLM and diffusion model with their task specific adapters separately. Due to the different styles of captions in different fashion datasets, we adopt the approach of instruction tuning to train the LLM so that it can generate captions of different styles. More details can be found in Appendix 3.3.
Phase 2:
In order to make UniFashion have the composed retrieval and generation abilities, we freeze the parameters of LLM and diffusion model, only fine-tune the multimodal encoder.
4.2 Datasets
We test the effectiveness of UniFashion by experimenting on different tasks including fashion image captioning, cross-modal retrieval, composed image retrieval and composed image generation.
For fashion image captioning tasks, we utilize the FashionGen (Zang et al., 2021) dataset. Additionally, to enhance our model’s capability in the CIR task, which involves the ability to retrieve captions for target images, we have annotated images from the training set of Fashion-IQ. Recognizing that manually annotating all the images would be time-consuming and resource-intensive, we draw inspiration from the success of recent MLLM models such as LLaVA in text-annotation tasks, and propose leveraging LLaVA 1.5 (13B) to semi-automatically annotate the dataset. More details can be found in Appendix C.
@@ 4.3 Evaluation Methods
Cross-modal Retrieval Evaluation.
We consider both image-to-text retrieval and text-to-image retrieval with random 100 protocols used by previous methods. 100 candidates are randomly sampled from the same category to construct a retrieval database. The goal is to locate the positive match depicting the same garment instance from these 100 same-category negative matches. We utilize Recall@K as the evaluation metric, which reflects the percentage of queries whose true target ranked within the top K candidates.
Fashion Image Captioning Evaluation.
For evaluating the performance of caption generation, we utilize BLEU-4, METEOR, ROUGE-L, and CIDEr as metrics.
Composed Fashion Image Retrieval Evaluation.
We compare our UniFashion with CIR methods and the FAME-ViL model of V + L that is oriented towards fashion in the original protocol used by Fashion-IQ. For this task, we also utilize Recall@K as the evaluation metric.
Composed Fashion Image Generation Evaluation.
We compare our UniFashion with try-on methods on VITON-HD dataset and fashion design works on MGD dataset. To evaluate the quality of image generation, we use the Frechet Inception Distance (FID) score to measure the divergence between two multivariate normal distributions and employ the CLIP Score (CLIP-S) provided in the TorchMetrics library to assess the adherence of the image to the textual conditioning input (for fashion design task).
@@ 4.4 Comparative Analysis of Baselines and Our Method
UniFashion exhibits superior performance across all datasets compared to baselines. Tab. 1 presents the evaluation results for each baseline and our models in FashionGen data sets for cross-modal retrieval. UniFashion outperforms most of the baseline models on both the text-to-image and image-to-text tasks. Following FAME-ViL, we also adopt a more challenging and practical protocol that conducts retrieval on the entire product set, which is in line with actual product retrieval scenarios. In Tab. 2, we performed a comparison between our UniFashion and other baselines on the FashionGen dataset for the image captioning task. By integrating the powerful generative ability of the LLM, our model performed significantly better than the traditional multimodal models in this task. In Tab. 4, we conducted a comparison between our UniFashion and CIR-specialist methods. Our findings are in line with those of Tab. 1.
After fine-tuning UniFashion on image generation/editing tasks with multimodal inputs, it exhibits outstanding performance. Tab. 3 evaluates the quality of the generated image of UniFashion in the VITON-HD unpaired setting. In order to verify that our model can achieve good results in a variety of modal inputs, we have conducted tests, respectively, on the traditional try-on task and the fashion design task proposed in MGD. For a fair evaluation with baselines, all the models are trained at a 512 × 384 resolution. To confirm the efficacy of our approach, we assess the realism using FID and KID score on all the tasks and using CLIP-S score for fashion design task. As can be seen, the proposed UniFashion model consistently outperforms competitors in terms of realism (i.e., FID and KID) and coherence with input modalities (i.e., CLIP-S), indicating that our method can better encode multimodal information. Meanwhile, although our model is slightly lower than StableVITON on the try-on task, this is because we froze the parameters of the diffusion model on the try-on task and only fine-tuned the Q-former part, but it can still achieve top2 results. The visual results can be found in Appendix E.
@@ 4.5 Ablation Study
UniFashion allows for more flexible execution of multimodal composed tasks. In Tab. 4, we also carry out ablation studies on different retrieval methods. Since UniFashion is capable of generating captions, for the CIR task, we initially utilize UniFashion to generate the captions of candidate images and then conduct the image retrieval task (denoted as UniFashion w/o cap) and the caption retrieval task (denoted as UniFashion w/o img). We find that our single-task variant has already achieved superior performance in the relevant field. Furthermore, due to the generative ability of our model, the pregenerated candidate library optimizes the model’s performance in this task. For specific implementation details, please refer to Appendix C.
We investigate the impact of different modules in UniFashion on various fashion tasks. In Tab. 5, we perform an ablation study on the proposed model architecture, with a focus on LLM and diffusion models. For comparison on the cross-modal retrieval task (CMR), we design the base model as directly fine-tuning BLIP2 without any new modules. The results indicate that the base model performs relatively well on this task and that the introduction of other modules does not lead to significant improvements. However, in the CIR task, the introduction of LLM and diffusion models as supervision can lead to significant improvements, especially when utilizing pregenerated captions by UniFashion to assist in retrieval, resulting in greater benefits. At the same time, we note that, after introducing the diffusion model, it may have some negative impact on the model’s image captioning ability, possibly due to the inherent alignment differences between LLM and the diffusion model.